OpenStudy (jchick):

1 Million Dollar Reward for Math answer: Fermat's Last Theorem - the idea that A^x + B^y = C^z. There could only be solutions to the equation when A, B, and C have a common numerical factor.

OpenStudy (jchick):

Not kidding there is a billionaire looking for the answer to this and he has offered 1 million dollars.

OpenStudy (anonymous):

IK! I saw this on ABC

OpenStudy (jchick):

A^x + B^y = C^z

OpenStudy (anonymous):

Idk understand but I'll say 1 and 2

OpenStudy (jchick):

I mean solve for A by simplifying both sides of the equation, then isolating the variable.

OpenStudy (jchick):

Take the logarithm of both sides of the equation to remove the variable from the exponent. x=ln(−By+Cz)ln(A)

OpenStudy (jchick):

to solve for x

OpenStudy (jchick):

Take the logarithm of both sides of the equation to remove the variable from the exponent. y=ln(−Ax+Cz)ln(B)

OpenStudy (jchick):

@mathmale

OpenStudy (jchick):

to solve for y

OpenStudy (jchick):

To solve for X Take the logarithm of both sides of the equation to remove the variable from the exponent. z=ln(Ax+By)ln(C)

OpenStudy (jchick):

Oops I meant Z

OpenStudy (jchick):

Solve for B by simplifying both sides of the equation, then isolating the variable. By=−Ax+Cz

OpenStudy (jchick):

Solve for C by simplifying both sides of the equation, then isolating the variable. Cz=Ax+By

OpenStudy (jchick):

If I am correct it cannot be solved.

OpenStudy (jchick):

given how exponentiation works even on the lower numbers you can quickly tell it won't work from powers of 2 and greater.

OpenStudy (jchick):

you can look at the powers of 2 to see this much. 3^2 = 9 which is 3*3 which is 3+3+3 4^2 = 16 which is 4*4 which is 4+4+4+4 see the problem if A B and C were all raised equally then the equation would work, but they aren't raised equally and the greater number you use for A B or C the greater the difference. it might work if you alter the powers

OpenStudy (jchick):

A^N + B^M = Z^O or perhaps A^N + B^N = Z^M where A = B

OpenStudy (thomas5267):

Fermat's last theorem has long been proven. The statement younsought to prove isn't even the statement in Fermat's last theorem.

OpenStudy (thomas5267):

*you sought

OpenStudy (anonymous):

fermat's last theorem: an + bn = cn

OpenStudy (anonymous):

or A^n+B^n=C^n

OpenStudy (thomas5267):

But he wrote Ax+By=Cz. Am I suppose to interpret that as A^x+B^y=C^z?

OpenStudy (jchick):

350-year-old mystery of Fermat's Last Theorem - the idea that A^x + B^y = C^z

OpenStudy (jchick):

When I wrote it I forgot to put the ^ symbols

OpenStudy (jchick):

And yes it does make a difference.

OpenStudy (thomas5267):

The statement should be: For integer n>2, a^n+b^n=c^n admits no solution.

OpenStudy (jchick):

This is the problem that the billionaire posted.

OpenStudy (thomas5267):

Proven by Andrew Wiles in 1995.

OpenStudy (thomas5267):

I am not messing with you. https://en.m.wikipedia.org/wiki/Fermat%27s_Last_Theorem

OpenStudy (jchick):

@thomas5267 then why is he offering a reward?

OpenStudy (jchick):

The margin note became known as Fermat’s Last Theorem,[17] as it was the last of Fermat’s asserted theorems to remain unproven

OpenStudy (thomas5267):

Probably the article claiming a billionaire offering a million dollar is a parody of some sort?

OpenStudy (jchick):

this is from that document or page you shared.

OpenStudy (thomas5267):

Second paragraph. "The first successful proof was released in 1994 by Andrew Wiles, and formally published in 1995, after 358 years of effort by mathematicians. "

OpenStudy (jchick):

However without this part proven, there was no actual proof of Fermat's Last Theorem. Wiles spent almost a year trying to repair his proof, initially by himself and then in collaboration with Richard Taylor, without success.[119] On 19 September 1994, on the verge of giving up, Wiles had a flash of insight that the proof could be saved by returning to his original Horizontal Iwasawa theory approach, which he had abandoned in favour of the Kolyvagin–Flach approach, this time strengthening it with expertise gained in Kolyvagin–Flach's approach.[120] On 24 October 1994, Wiles submitted two manuscripts, "Modular elliptic curves and Fermat's Last Theorem"[121] and "Ring theoretic properties of certain Hecke algebras",[122] the second of which was co-authored with Taylor and proved that certain conditions were met that were needed to justify the corrected step in the main paper. The two papers were vetted and published as the entirety of the May 1995 issue of the Annals of Mathematics. These papers established the modularity theorem for semistable elliptic curves, the last step in proving Fermat's Last Theorem, 358 years after it was conjectured.

OpenStudy (jchick):

How?

OpenStudy (anonymous):

Jk I wish I knew.

OpenStudy (jchick):

lol

OpenStudy (anonymous):

It's really complicated, you need to have read "Annals of Mathematics" to come even close

OpenStudy (thomas5267):

Fermat's last theorem is proven. What the billionare wants is the Beal conjecture. The article is so poorly written no wonder you are confused.

OpenStudy (jchick):

Ok I see it now. I misunderstood it then because ABC wrote it lol anyways if they are trying to solve the beal conjecture let me find that.

OpenStudy (jchick):

Oh yes I thought that they were saying in the article that he was wrong when he supposedly proved it.

OpenStudy (jchick):

But yes as you have said it was solved by Andrew Wiley in 1994

jimthompson5910 (jim_thompson5910):

Fermat's Last Theorem is this \[\Large x^n + y^n = z^n\] has no non-trivial integer solutions when n > 2 source: http://www-history.mcs.st-and.ac.uk/HistTopics/Fermat's_last_theorem.html

jimthompson5910 (jim_thompson5910):

I don't know where that page is getting Ax + By = Cz

jimthompson5910 (jim_thompson5910):

oh, maybe they mean to say A^x + B^y = C^z

OpenStudy (jchick):

It is actually A^x + B^y = C^z

OpenStudy (jchick):

Beal's Conjecture A generalization of Fermat's last theorem which states that if a^x+b^y=c^z, where a, b, c, x, y, and z are any positive integers with x,y,z>2, then a, b, and c have a common factor. The conjecture was announced in Mauldin (1997), and a cash prize of $1000000 has been offered for its proof or a counterexample

jimthompson5910 (jim_thompson5910):

these two pages state that the theorem has been finally proved http://mathworld.wolfram.com/FermatsLastTheorem.html http://mathforum.org/dr.math/faq/faq.fermat.html so I think there's something off here

OpenStudy (jchick):

Yes we have determined that it is Beal's conjecture

OpenStudy (thomas5267):

I wouldn't be surprised if someone is working on Beal conjecture now. After all Andrew Wiles worked alone for 6 years before coming up with a proof. Just to clarify, Fermat's last theorem was proven by Andrew Wiles in 1995. The billionaire Andy Beal want the proof of a different conjecture named Beal conjecture and will give a million dollar to the person who proved it. Beal conjecture is a generalisation of Fermat's last theorem.

OpenStudy (jchick):

1994 but yes.

OpenStudy (jchick):

Also @thomas5267 I wouldn't recommend using Wikipedia as a source.

jimthompson5910 (jim_thompson5910):

ah I see now, nvm what I posted

OpenStudy (jchick):

But since there are so many other sources then it could very well be plausible.

OpenStudy (thomas5267):

Wikipedia is more than alright when you are engaging in an informal discussion. When you are writing any formal essay, Wikipedia is a big no-no.

OpenStudy (jchick):

Yes I was just simply pointing out that Wikipedia is not necessarily a very good source. Yes for these purposes it is alright.

OpenStudy (thomas5267):

I won't use Wikipedia if my grades depends on it. I do use Wikipedia as a starting point and for the references in Wikipedia.

OpenStudy (jchick):

So since everyone on OpenStudy is looking at this any ideas?

OpenStudy (jchick):

lol jk

OpenStudy (jchick):

Sorry but I truly need to leave now I will talk to and maybe help many of you later after I am done with tutoring tonight.

jimthompson5910 (jim_thompson5910):

this person/page invested a lot of time into trying to find counterexamples http://norvig.com/beal.html but didn't find any. So it means that if there are counterexamples, then they are so large that they won't be found any time soon. If there are no counterexamples, then the actual proof itself is going to be nearly impossible unless you're a very gifted mathematician

jimthompson5910 (jim_thompson5910):

so there's a reason why there's a lot of money put up for this problem